OpenStudy (anonymous):

What is the radius of a circle in which a 30° arc is 2 inches long?

OpenStudy (anonymous):

2pi*

OpenStudy (anonymous):

2 pi r 30/360 =2

OpenStudy (anonymous):

2 pi r 1/12= 2 pi r 2/12= 2

OpenStudy (anonymous):

2*(12/2)=12 pi r= 2 r= 2/pi

OpenStudy (anonymous):

Thank you!